OpenStudy (anonymous):

Menthol is a crystalline substance with a peppermint taste and odor. When 1.45 g of menthol is dissolved in 25.0 g of cyclohexane, the freezing point of the solution is lowered by 7.73 °C. The freezing point and Kf constant for cyclohexane can be found here. Calculate the molar mass of menthol.

OpenStudy (aaronq):

use the formula: \(\Delta T=i*m*K_{f}\) find molality, then expand the formula ; m=\(\dfrac{n_{solute}}{kg_{solvent}}=\dfrac{\dfrac{m_{menthol}}{M_{menthol}}}{kg_{solvent}}\) More on this: http://openstudy.com/study#/updates/521ecb3ae4b0750826e0c362